# Redshift Dipoles from Non-Geodesic Observer Congruences in Covariant Cosmology

### 21 Mar 2026 v1

@@ Abstract
In an inhomogeneous universe, the Hubble frame used to describe the cosmic expansion does not, in general, coincide with a geodesic matter flow. In this work, it is shown that within a fully covariant framework, a non-geodesic observer congruence introduces an additional contribution to the relation between affine parameter and observed redshift, proportional to the line-of-sight projection of the observer four-acceleration. This induces a dipolar modulation of the redshift itself, which propagates to any observable expressed in redshift space.
@@
Unlike the standard kinematic dipole associated with a global Lorentz boost, this contribution is sourced by the kinematics of the observer congruence along the past light cone and can exhibit a non-trivial redshift dependence. These results provide a direct framework to identify such signatures in cosmological data, with potential implications for the interpretation of large-scale dipoles and redshift-based inference.
1 Introduction
@@
Cosmological observables are defined on the past light–cone of the observer, reflecting the causal structure of spacetime [5, 6]. Among them, redshift plays a central role, encoding the relation between spacetime dynamics and observational data. In standard cosmology, it is interpreted with respect to the comoving geodesic congruence, for which it reduces to the familiar relation $1+z=a_{0}/a$.
@@
However, this identification implicitly assumes a particular choice of observer congruence. In realistic, inhomogeneous spacetimes, the congruence used to define cosmological observables—such as those associated with the Hubble flow or the CMB rest frame—need not coincide with the geodesic matter frame. In general, different congruences are related by spacetime-dependent local boosts, under which geodesicity is not preserved. As a result, an observer congruence may possess a non-vanishing four-acceleration even in a spacetime sourced by pressureless matter. This raises a fundamental question: how does the choice of observer congruence affect the definition and interpretation of cosmological observables?
@@
In this work, the problem is addressed within a fully covariant framework [5, 12]. The optical focusing equation governing the angular diameter distance is purely geometrical, while the relation between affine parameter and observed redshift depends explicitly on the kinematics of the observer congruence. In particular, a non-geodesic congruence introduces additional contributions to the redshift propagation equation, leading to a dipolar modulation proportional to the projection of the four-acceleration along the line of sight. This result has immediate observational consequences. Since cosmological distances are expressed as functions of redshift, any dipolar modulation in the redshift propagates to all redshift-space observables. Importantly, this dipole originates from the local kinematics of the observer congruence and is not, in general, reducible to a single global Lorentz boost.
@@
An interesting observational point is that accelerated reference frames are not merely a theoretical possibility. The acceleration of the Solar System barycentre produces the well-known secular aberration drift [11, 2], a dipolar proper-motion pattern that has been measured with very long baseline interferometry and more recently constrained with Gaia astrometry. This shows that observer acceleration can leave directly observable anisotropic signatures in precision cosmological measurements, and it motivates considering whether analogous kinematic effects may also enter the definition of redshift when the relevant cosmological frame is not geodesic. More broadly, several large-scale dipolar or directional anomalies have been discussed in the literature, including the low-multipole alignments in the CMB and the unexpectedly large dipole reported in quasar number counts (see e.g. [7, 10, 3]). While such observations need not be explained by the mechanism proposed here, they illustrate that the interpretation of cosmic dipoles remains open and deserves careful scrutiny. In that context, the possibility that part of the observed anisotropy may arise from the kinematics of a non-geodesic observer congruence is worth considering, precisely because it does not rely on new degrees of freedom or modifications of gravity, but follows from the covariant relativistic description itself.
@@
The structure of this paper is as follows. The Sachs optical equation for the angular diameter distance is first derived, and its purely geometrical character is emphasized. The propagation of redshift in both geodesic and non-geodesic observer congruences is then analyzed, showing that a dipolar modulation naturally arises in the latter case. Order-of-magnitude estimates for the expected amplitude of this effect are presented, and an operational observable to isolate the dipolar component in redshift data is introduced. The paper concludes with a discussion of the implications for cosmological inference and possible observational tests.
2 The angular diameter distance from the Raychaudhuri equation
@@
$k^{\beta}\nabla_{\beta}k^{\alpha}=0,\qquad k^{\alpha}k_{\alpha}=0.$ (2.1)
@@
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (2.2)
@@
$\sigma_{N}^{2}\equiv\frac{1}{2}\,\sigma^{(N)}_{\alpha\beta}\sigma_{(N)}^{\alpha\beta}\geq 0,\qquad\omega_{N}^{2}\equiv\frac{1}{2}\,\omega^{(N)}_{\alpha\beta}\omega_{(N)}^{\alpha\beta}\geq 0,$ (2.3)
@@
$\frac{d\theta_{N}}{d\lambda}=-\frac{1}{2}\theta_{N}^{2}-2\sigma_{N}^{2}+2\omega_{N}^{2}-R_{\alpha\beta}k^{\alpha}k^{\beta}.$ (2.4)
@@
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}\left(R_{\alpha\beta}k^{\alpha}k^{\beta}+2\sigma_{N}^{2}\right)\sqrt{A}.$ (2.6)
@@
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\sqrt{A}.$ (2.7)
@@
$d_{A}^{2}\equiv\frac{A}{\Omega},$ (2.8)
@@
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A}.$ (2.9)
Equation (2.9) corresponds to the shear-free limit of the Sachs optical equation [9], valid for hypersurface-orthogonal null congruences with vanishing vorticity. In this case, the evolution of the angular diameter distance is governed solely by the Ricci focusing term. The initial conditions at the observer (located at $\lambda=0$) are
$d_{A}(0)=0,\qquad\left.\frac{dd_{A}}{d\lambda}\right|_{0}=1,$ (2.10)
@@
2.1 Matter source term and FLRW universe
@@
$R_{\alpha\beta}=8\pi G\left(T_{\alpha\beta}-\frac{1}{2}Tg_{\alpha\beta}\right)+\Lambda g_{\alpha\beta},$ (2.11)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G\,T_{\alpha\beta}k^{\alpha}k^{\beta}.$ (2.12)
@@
$T_{\alpha\beta}=(\rho+p)u_{\alpha}u_{\beta}+pg_{\alpha\beta}.$ (2.13)
@@
$T_{\alpha\beta}k^{\alpha}k^{\beta}=(\rho+p)(u\cdot k)^{2}.$ (2.14)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)(u\cdot k)^{2}.$ (2.15)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)E^{2}.$ (2.16)
As a consistency check, specialising to an FLRW spacetime where the null congruence associated with radial light propagation is shear-free and the metric takes its standard form, the solution of Eq. (2.9) is (see Appendix B)
$d_{A}=a(\eta)\,S_{K}\!\big(\chi\big).$ (2.17)
@@
3 Redshift definition and the geodesic observer congruence
@@
$1+z_{obs}\equiv\frac{(u\cdot k)_{\rm em}}{(u\cdot k)_{\rm obs}},\qquad E\equiv-(u\cdot k),$ (3.1)
@@
$\frac{dE}{d\lambda}=-\,k^{\alpha}\nabla_{\alpha}(u_{\beta}k^{\beta})=-\,k^{\alpha}k^{\beta}\nabla_{\alpha}u_{\beta},$ (3.2)
@@
$\nabla_{\alpha}u_{\beta}=\frac{1}{3}\theta\,h_{\alpha\beta}+\sigma_{\alpha\beta}+\omega_{\alpha\beta}-u_{\alpha}\mathcal{A}_{\beta},\qquad h_{\alpha\beta}=g_{\alpha\beta}+u_{\alpha}u_{\beta}.$ (3.3)
@@
$k^{\alpha}=E\,(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,$ (3.4)
@@
Substituting (3.3) and (3.4) into (3.2), using $k^{\alpha}k_{\alpha}=0$ and $h_{\alpha\beta}u^{\beta}=0$, one obtains the exact kinematical propagation law:
$\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta+\sigma_{\alpha\beta}e^{\alpha}e^{\beta}-\mathcal{A}_{\alpha}e^{\alpha}\right).$ (3.5)
@@
$\mathcal{A}^{\alpha}\equiv u^{\beta}\nabla_{\beta}u^{\alpha}=0.$ (3.6)
@@
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,\frac{1}{3}E\theta.}$ (3.7)
Equation (3.7) reduces to the standard FLRW redshift law. Indeed, defining the expansion scalar and local Hubble rate of the observer congruence by
$H\equiv\frac{\theta}{3},$ (3.8)
@@
$\frac{1}{E}\frac{dE}{d\lambda}=-EH.$ (3.9)
@@
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH.$ (3.10)
@@
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha},\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (3.11)
@@
$\frac{da}{d\lambda}=k^{\alpha}\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}k^{\alpha}=\dot{a}\,E=aHE.$ (3.12)
@@
$\frac{1}{a}\frac{da}{d\lambda}=EH.$ (3.13)
Comparing (3.10) and (3.13), we obtain
$\frac{d\ln(1+z)}{d\lambda}=-\frac{d\ln a}{d\lambda},$ (3.14)
@@
$1+z_{obs}=\frac{a_{0}}{a},$ (3.15)
where $a$ is the scale factor evaluated at the source position. Thus, Eq. (3.5) provides the exact covariant propagation law whose geodesic, shear-free FLRW limit reproduces the standard cosmological redshift relation.
4 Non-geodesic observer congruence and the dipole modulation
The Sachs equation (2.9) for $d_{A}$ is purely geometric and does not require the observer congruence to be geodesic. In the FLRW derivation above, the assumption of geodesicity entered only when we related the affine parameter (or conformal time) to the redshift parameter $z$ defined along the null geodesic. We now relax this assumption and make explicit where and how the modification appears.
In the previous FLRW derivation we implicitly took $u^{\alpha}$ to be the comoving geodesic congruence. We now relax (3.6) and allow
$\mathcal{A}^{\alpha}\neq 0.$ (4.1)
The impact of (4.1) is entirely contained in the propagation of $E$ along the null ray. Neglecting the shear of the observer congruence for simplicity, the propagation law reduces to
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta-\mathcal{A}_{\alpha}e^{\alpha}\right).}$ (4.2)
Note that the $\mathcal{A}_{\alpha}e^{\alpha}$ term is direction-dependent. Equation (4.2) can be written as
$\frac{1}{E}\frac{dE}{d\lambda}=-EH+E\,\mathcal{A}_{\alpha}e^{\alpha},$ (4.3)

### 14 Apr 2026 v2

@@ Abstract
In an inhomogeneous universe, the identification of the Hubble frame with a geodesic matter flow is not guaranteed beyond the idealized FLRW limit, particularly once structure formation leads to a non-trivial distribution of velocities and gravitational fields. Within a fully covariant framework, we show that a non-geodesic observer congruence introduces an additional contribution to the propagation of redshift along the past light cone, proportional to the line-of-sight projection of the observer four-acceleration. This generates a dipolar modulation in the redshift itself, which propagates to any observable defined in redshift space.
@@
Unlike the standard kinematic dipole associated with a global Lorentz boost, this contribution arises from the kinematics of the observer congruence and depends on its evolution along the light cone, leading to a non-trivial redshift dependence. These results provide a framework to explore potential signatures of non-geodesic kinematics in cosmological data, with possible implications for the interpretation of large-scale dipoles and redshift-based inference.
I Introduction
@@
Cosmological observables are defined on the past light–cone of the observer, reflecting the causal structure of spacetime [5, 7]. Among them, redshift plays a central role, encoding the relation between spacetime dynamics and observational data. In standard cosmology, it is interpreted with respect to the comoving geodesic congruence, for which it reduces to the familiar relation $1+z=a_{0}/a$. However, this interpretation relies on a specific identification between the observer congruence and the underlying matter flow.
In realistic, inhomogeneous spacetimes, the congruence used to define cosmological observables—such as those associated with the Hubble flow or the CMB rest frame—is typically defined through local kinematic corrections and need not correspond to a globally defined geodesic matter flow along the past light cone. This mismatch is expected to become more relevant at low redshift, where structure formation induces significant inhomogeneities, peculiar velocity fields, and anisotropic gravitational environments.
@@
In general, different congruences are related by spacetime-dependent local boosts, under which geodesicity is not preserved. As a result, an observer congruence may possess a non-vanishing four-acceleration even in a spacetime sourced by pressureless matter. Here, the four-acceleration does not describe a force acting on matter, but rather the non-geodesic nature of the observer congruence used to define redshift. This raises a fundamental question: how does the choice of observer congruence, and in particular its possible non-geodesic character, affect the definition and interpretation of cosmological observables? It is important to emphasize that the present analysis does not question the standard practice of applying local kinematic corrections to define observational frames such as the CMB rest frame. Rather, it highlights that such corrections do not uniquely define a geodesic observer congruence along the past light-cone, and therefore need not correspond to a physically global Hubble flow in an inhomogeneous universe. Therefore, the observational relevance of the four-acceleration arises because redshift is operationally defined with respect to this congruence.
@@
In this work, the problem is addressed within a fully covariant framework [5, 16]. The optical focusing equation governing the angular diameter distance is purely geometrical, while the relation between affine parameter and observed redshift depends explicitly on the kinematics of the observer congruence. In particular, a non-geodesic congruence introduces additional contributions to the redshift propagation equation along the past light cone, leading to a dipolar modulation proportional to the projection of the four-acceleration along the line of sight. This result has immediate observational consequences. Since cosmological distances are expressed as functions of redshift, any dipolar modulation in the redshift propagates to all redshift-space observables. Importantly, this dipole originates from the local kinematics of the observer congruence and is not, in general, reducible to a single global Lorentz boost.
@@
An interesting observational point is that accelerated reference frames are not merely a theoretical possibility. The acceleration of the Solar System barycentre produces the well-known secular aberration drift [15, 3], a dipolar proper-motion pattern that has been measured with very long baseline interferometry and more recently constrained with Gaia astrometry. This illustrates that observer acceleration can leave directly observable anisotropic signatures in precision cosmological measurements, and it motivates considering whether analogous kinematic effects may also enter the definition of redshift when the relevant cosmological frame is not geodesic. More broadly, several large-scale dipolar or directional anomalies have been discussed in the literature, including the low-multipole alignments in the CMB and the unexpectedly large dipole reported in quasar number counts (see e.g. [10, 13, 4]). While such observations need not be explained by the mechanism proposed here, they illustrate that the interpretation of cosmic dipoles remains open and deserves careful scrutiny. In that context, the possibility that part of the observed anisotropy may arise from the kinematics of a non-geodesic observer congruence is worth considering, precisely because it follows from the covariant relativistic description itself, without invoking additional degrees of freedom or modifications of gravity.
@@
The structure of this paper is as follows. The Sachs optical equation for the angular diameter distance is first derived, and its purely geometrical character is emphasized. The propagation of redshift in both geodesic and non-geodesic observer congruences is then analyzed, showing that a dipolar modulation naturally arises in the latter case. Order-of-magnitude estimates for the expected amplitude of this effect are presented, and an operational observable designed to isolate the dipolar component in redshift data is introduced. The paper concludes with a discussion of the implications for cosmological inference and possible observational tests.
II The angular diameter distance from the Raychaudhuri equation
@@
$k^{\beta}\nabla_{\beta}k^{\alpha}=0,\qquad k^{\alpha}k_{\alpha}=0.$ (1)
@@
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (2)
@@
$\sigma_{N}^{2}\equiv\frac{1}{2}\,\sigma^{(N)}_{\alpha\beta}\sigma_{(N)}^{\alpha\beta}\geq 0,\qquad\omega_{N}^{2}\equiv\frac{1}{2}\,\omega^{(N)}_{\alpha\beta}\omega_{(N)}^{\alpha\beta}\geq 0,$ (3)
@@
$\frac{d\theta_{N}}{d\lambda}=-\frac{1}{2}\theta_{N}^{2}-2\sigma_{N}^{2}+2\omega_{N}^{2}-R_{\alpha\beta}k^{\alpha}k^{\beta}.$ (4)
@@
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}\left(R_{\alpha\beta}k^{\alpha}k^{\beta}+2\sigma_{N}^{2}\right)\sqrt{A}.$ (6)
@@
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\sqrt{A}.$ (7)
@@
$d_{A}^{2}\equiv\frac{A}{\Omega},$ (8)
@@
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A}.$ (9)
Equation (9) corresponds to the shear-free limit of the Sachs optical equation [12], valid for hypersurface-orthogonal null congruences with vanishing vorticity. In this case, the evolution of the angular diameter distance is governed solely by the Ricci focusing term. The initial conditions at the observer (located at $\lambda=0$) are
$d_{A}(0)=0,\qquad\left.\frac{dd_{A}}{d\lambda}\right|_{0}=1,$ (10)
@@
II.1 Matter source term and FLRW universe
@@
$R_{\alpha\beta}=8\pi G\left(T_{\alpha\beta}-\frac{1}{2}Tg_{\alpha\beta}\right)+\Lambda g_{\alpha\beta},$ (11)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G\,T_{\alpha\beta}k^{\alpha}k^{\beta}.$ (12)
@@
$T_{\alpha\beta}=(\rho+p)u_{\alpha}u_{\beta}+pg_{\alpha\beta}.$ (13)
@@
$T_{\alpha\beta}k^{\alpha}k^{\beta}=(\rho+p)(u\cdot k)^{2}.$ (14)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)(u\cdot k)^{2}.$ (15)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)E^{2}.$ (16)
As a consistency check, specialising to an FLRW spacetime where the null congruence associated with radial light propagation is shear-free and the metric takes its standard form, the solution of Eq. (9) is (see Appendix B)
$d_{A}=a(\eta)\,S_{K}\!\big(\chi\big).$ (17)
@@
III Redshift definition and the geodesic observer congruence
@@
$1+z_{obs}\equiv\frac{(u\cdot k)_{\rm em}}{(u\cdot k)_{\rm obs}},\qquad E\equiv-(u\cdot k),$ (18)
@@
$\frac{dE}{d\lambda}=-\,k^{\alpha}\nabla_{\alpha}(u_{\beta}k^{\beta})=-\,k^{\alpha}k^{\beta}\nabla_{\alpha}u_{\beta},$ (19)
@@
$\nabla_{\alpha}u_{\beta}=\frac{1}{3}\theta\,h_{\alpha\beta}+\sigma_{\alpha\beta}+\omega_{\alpha\beta}-u_{\alpha}\mathcal{A}_{\beta},\qquad h_{\alpha\beta}=g_{\alpha\beta}+u_{\alpha}u_{\beta}.$ (20)
@@
$k^{\alpha}=E\,(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,$ (21)
@@
Substituting (20) and (21) into (19), using $k^{\alpha}k_{\alpha}=0$ and $h_{\alpha\beta}u^{\beta}=0$, one obtains the exact kinematical propagation law:
$\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta+\sigma_{\alpha\beta}e^{\alpha}e^{\beta}-\mathcal{A}_{\alpha}e^{\alpha}\right).$ (22)
@@
$\mathcal{A}^{\alpha}\equiv u^{\beta}\nabla_{\beta}u^{\alpha}=0.$ (23)
@@
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,\frac{1}{3}E\theta.}$ (24)
Equation (24) reduces to the standard FLRW redshift law. Indeed, defining the expansion scalar and local Hubble rate of the observer congruence by
$H\equiv\frac{\theta}{3},$ (25)
@@
$\frac{1}{E}\frac{dE}{d\lambda}=-EH.$ (26)
@@
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH.$ (27)
@@
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha},\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (28)
@@
$\frac{da}{d\lambda}=k^{\alpha}\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}k^{\alpha}=\dot{a}\,E=aHE.$ (29)
@@
$\frac{1}{a}\frac{da}{d\lambda}=EH.$ (30)
Comparing (27) and (30), we obtain
$\frac{d\ln(1+z)}{d\lambda}=-\frac{d\ln a}{d\lambda},$ (31)
@@
$1+z_{obs}=\frac{a_{0}}{a},$ (32)
where $a$ is the scale factor evaluated at the source position. Thus, Eq. (22) provides the exact covariant propagation law whose geodesic, shear-free FLRW limit reproduces the standard cosmological redshift relation.
IV Non-geodesic observer congruence and the dipole modulation
The Sachs equation (9) for $d_{A}$ is purely geometrical and does not require the observer congruence to be geodesic. In the FLRW derivation above, the assumption of geodesicity entered only when relating the affine parameter (or conformal time) to the redshift parameter $z$ defined along the null geodesic. We now relax this assumption and make explicit where and how the modification appears.
In the previous FLRW derivation we implicitly took $u^{\alpha}$ to be the comoving geodesic congruence. We now relax (23) and allow
$\mathcal{A}^{\alpha}\neq 0.$ (33)
The impact of (33) is entirely contained in the propagation of $E$ along the null ray. In full generality, the propagation law is
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta+\sigma_{\alpha\beta}e^{\alpha}e^{\beta}-\mathcal{A}_{\alpha}e^{\alpha}\right).}$ (34)
The different kinematical contributions have a clear angular structure: the expansion scalar $\theta$ contributes to the monopole, the projection $\mathcal{A}_{\alpha}e^{\alpha}$ generates a dipolar term, and the shear contribution $\sigma_{\alpha\beta}e^{\alpha}e^{\beta}$ induces a quadrupolar modulation.
Equation (34) can be written as
$\frac{1}{E}\frac{dE}{d\lambda}=-EH-E\,\sigma_{\alpha\beta}e^{\alpha}e^{\beta}+E\,\mathcal{A}_{\alpha}e^{\alpha},$ (35)